# On estimates of trace-norm distance between quantum Gaussian states

### v2

@@ Abstract
In the paper [1] estimates for the trace-norm distance between two quantum Gaussian states in terms of the mean vectors and covariance matrices were derived and used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states. In the present paper we obtain different estimates; our proof is based on a fidelity-like quantity which we call states overlap, and is more straightforward leading to estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. These derivations are not aimed to replace the useful inequalities from [1]; they just show an alternative approach to the problem leading to different results. The problem studied in this paper can be considered as a noncommutative analog of estimation of the total variance distance between Gaussian probability distributions in the classical probability theory.
@@ 1 Introduction and results
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq 2\sqrt{8\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
The proofs are presented in sec. 3. In the concluding sec. 4 we will compare our estimates with those obtained in [1].
@@ 2 The algebra of CCR
defining the symplectic form $\Delta(z,z^{\prime})=z^{t}\Delta z^{\prime}.$ In what follows the Hilbert space $\mathcal{H}$ is the space of an irreducible representation $z\rightarrow W(z);\,z\in Z,$ of the Weyl canonical commutation relations
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
where $\rho_{1},\rho_{2}$ are density operators. If the corresponding states are pure then the second inequality turns into equality. The proof given in [7] is partly based on estimates from [10].
@@
where $T=\hat{\alpha}^{-1}\hat{\beta}$ is positive operator in the Hilbert space $Z_{\hat{\alpha}}.$ Also $T-2+T^{-1}$ is positive operator in $Z_{\hat{\alpha}},$ therefore
@@
$=$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\ln\left(I+\frac{T-2+T^{-1}}{4}\right)\right]$
$\geq$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\left(\frac{T-2+T^{-1}}{4}\right)\right],$
where the last inequality follows from $\ln\left(1+x\right)\leq x$ for $x\geq 0.$
Then from (29) it follows
$1-\left(\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}\right)^{2}$
$\leq$ $m^{t}\sigma(\hat{\alpha},\hat{\beta})^{-1}m+\frac{1}{8}\mathrm{Tr}\left(T-2+T^{-1}\right)$
$=$ $2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{8}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$
@@
Let us evaluate the expression $\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$
From (18)
@@ 3.1 Pure states
In the case of pure state the relation (11) implies $\left(2\Delta^{-1}\alpha\right)^{2}=-I_{2s},$ that is $\Upsilon_{\alpha}=0$ and $\hat{\alpha}=\alpha.$ Similarly $\hat{\beta}=\beta$ and $\Upsilon_{\beta}=0$. Then (33) with (30) imply
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{2m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{2}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{8\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\frac{1}{2}\left\|\alpha-\beta\right\|_{2}^{2}},$ (35)
@@ 4 Discussion
In the paper [1] also the lower bounds in terms of first and second moments were derived for the trace-norm distance of two Gaussian states (theorem 5). It would be interesting to investigate whether the first inequality in (26) could be used for a similar purpose.

### v3

@@ Abstract
In the paper [1] estimates for the trace-norm distance between two quantum Gaussian states of CCR in terms of the mean vectors and covariance matrices were derived and used to evaluate the number of elements in the $\varepsilon-$net in the set of energy-constrained Gaussian states. In the present paper we obtain different estimates based on a fidelity-like quantity which we call states overlap [2]. Our proof is more direct leading to estimates which are sometimes even more stringent, especially in the cases of pure or gauge-invariant states. They do not depend on number of modes and hence can be extended to the case of bosonic field with infinite number of modes. These derivations are not aimed to replace the useful inequalities from [1]; they just show an alternative approach to the problem leading to different results. In the Appendix we briefly recall our results from [3] concerning estimates of the overlap for general fermionic Gaussian states of CAR.
The problems studied in this paper can be considered as a noncommutative analog of estimation of the total variance distance between Gaussian probability distributions in the classical probability theory.
@@ 1 Introduction and results
$\left\|\rho_{m_{1},\alpha}-\rho_{{}_{m_{2},\beta}}\right\|_{1}\leq\sqrt{16\min\left(a,b\right)\left\|m_{1}-m_{2}\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (1)
@@
The proofs are presented in sec. 3. In the concluding sec. 4 we compare our estimates with those obtained in [1]. In the Appendix we briefly recall our results from [3] concerning estimates of overlap for general fermionic Gaussian states of CAR.
@@ 2 The algebra of CCR
defining the symplectic form $\Delta(z,z^{\prime})=z^{t}\Delta z^{\prime}.$ In what follows the Hilbert space $\mathcal{H}$ is the space of an irreducible representation $z\rightarrow W(z);\,z\in Z,$ of the Weyl canonical commutation relations (CCR)
@@ A Gaussian state $\rho_{m,\alpha}$ is determined by its quantum characteristic function
where $\rho_{1},\rho_{2}$ are density operators. The proof given in [2] is partly based on estimates from [11].
@@
where $T=\hat{\alpha}^{-1}\hat{\beta}$ is positive operator in the Hilbert space $Z_{\hat{\alpha}}.$ Also
$d(\alpha,\beta)=\frac{1}{4}\left(T-2+T^{-1}\right)$
is positive operator in $Z_{\hat{\alpha}},$ therefore
@@
$=$ $\exp\left[-\frac{1}{4}\mathrm{Tr}\ln\left(I+d(\alpha,\beta)\right)\right].$
By using the inequalities $\frac{\ln(1+c)}{c}x\leq\ln\left(1+x\right)\leq x$ for $0\leq x\leq c,$ we obtain
$\exp\left[-\frac{1}{4}\mathrm{Tr\,}d(\alpha,\beta)-m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right]\leq\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}$ (30)
$\leq$ $\exp\left[-\frac{1}{4}\frac{\ln(1+\left\|d(\alpha,\beta)\right\|)}{\left\|d(\alpha,\beta)\right\|}\mathrm{Tr\,}d(\alpha,\beta)-m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m\right].$ (31)
The first inequality implies
$1-\left(\mathrm{Tr}\sqrt{\rho_{\alpha}}\sqrt{\rho_{m,\beta}}\right)^{2}\leq 2m^{t}\left(\hat{\alpha}+\hat{\beta}\right)^{-1}m+\frac{1}{2}\mathrm{Tr\,}d(\alpha,\beta).$
But
$\mathrm{Tr\,}d(\alpha,\beta)=\frac{1}{4}\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$
@@
Let us evaluate the expression $\mathrm{Tr}\left(\hat{\alpha}-\hat{\beta}\right)\left(\hat{\beta}^{-1}-\hat{\alpha}^{-1}\right).$ From (18)
@@ 3.1 Pure states
In the case of general pure states there is exact expression (see e.g. Lemma 10.9 in [10]) which can be rewritten in the form
$\left\|\rho_{1}-\rho_{2}\right\|_{1}=2\sqrt{1-\mathrm{Tr}\rho_{1}\rho_{2}}=2\sqrt{1-\mathrm{Tr}\sqrt{\rho_{1}}\sqrt{\rho_{2}}}..$ (36)
Also for Gaussian pure states the relation (11) implies $\left(2\Delta^{-1}\alpha\right)^{2}=-I_{2s},$ that is $\Upsilon_{\alpha}=0$ and $\hat{\alpha}=\alpha.$ Similarly $\hat{\beta}=\beta$ and $\Upsilon_{\beta}=0$. Then (35) with (30) imply
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq 2\sqrt{m^{t}\left(\alpha+\beta\right)^{-1}m+\frac{1}{4}\mathrm{Tr}\left(\alpha-\beta\right)^{2}}.$
@@
$\left\|\rho_{\alpha}-\rho_{{}_{m,\beta}}\right\|_{1}\leq\sqrt{16\min\left(a,b\right)\left\|m\right\|_{2}^{2}+\left\|\alpha-\beta\right\|_{2}^{2}},$ (38)
@@ 4 Discussion
In the paper [1] also the lower bounds in terms of first and second moments were derived for the trace-norm distance of two Gaussian states (theorem 5). It would be interesting to investigate whether the second inequality in (31) could be used for a similar purpose.
5 Appendix: the case of fermionic Gaussian states of CAR
Here we briefly recall some results from our paper [3] where the overlap was computed for general fermionic Gaussian states of canonical anticommutation relations (CAR).
Let $H$ be a real even- or infinite-dimensional Hilbert space with inner product $(f,g)$. Let $\mathfrak{A}(H)$ the C-algebra of CAR generated by the elements $B(f),\,f\in H,$ where $f\rightarrow B(f)$ is a linear map and
$B(f)B(g)+B(g)B(f)=2(f,g)I.$
A state $\rho\left(\cdot\right)$ on $\mathfrak{A}(H)$ is called quasifree (also fermionic Gaussian) if it is even and
$\rho\left(B(f_{1})\dots B(f_{2n})\right)=\sum\left(-1\right)^{p}\rho\left(B(f_{i_{1}})B(f_{j_{1}})\right)\dots\rho\left(B(f_{i_{n}})B(f_{j_{n}})\right),$ (48)
[…]
$\rho\left(B(f)B(g)\right)=(f,g)+i(Af,g).$ (49)
The operator $A$ satisfies
$A^{\ast}=-A,\ \left\|A\right\|\leq 1,$ (50)
and it is called covariance operator of the state $\rho\left(\cdot\right).$ Conversely, any such operator defines a quasifree state according to (49), (48) which is denoted $\rho_{A}\left(\cdot\right).$ We denote by $\rho_{A}$ the density operator of the state $\rho_{A}(\cdot).$
One has $0\leq A^{\ast}A=-A^{2}\leq I.$ The operator $I+A^{2}$ is symmetric, $0\leq I+A^{2}\leq I,$ hence $\sqrt{I+A^{2}}$ is defined and $0\leq\sqrt{I+A^{2}}\leq I.$ The state $\rho_{A}$ is pure iff $I+A^{2}=0.$ For the proof of the following results see [3].
Lemma 1. Let $A,B$ be operators on $H$ satisfying (50), then the overlap of $\rho_{A},\rho_{B}$ is
$\mathrm{Tr}\sqrt{\rho_{A}}\sqrt{\rho_{B}}=\sqrt[4]{\det S^{\ast}S}$ (51)
where
$S$ $=$ $\frac{1}{2}[\left(I+\sqrt{I+B^{2}}\right)^{1/2}\left(I+\sqrt{I+A^{2}}\right)^{1/2}$
-- $B\left(I+\sqrt{I+B^{2}}\right)^{-1/2}A\left(I+\sqrt{I+A^{2}}\right)^{-1/2}]$
Lemma 2. $0\leq S^{\ast}S\leq I.$
Introducing $s(A,B)=\mathrm{Tr}\left(I-S^{\ast}S\right)$ we have the following estimates for the overlap:
$4]{1-s(A,B)}\leq\mathrm{Tr}\sqrt{\rho_{A}}\sqrt{\rho_{B}}\leq\exp\left(-\frac{s(A,B)}{4}\right),$ (52)
provided $s(A,B)\leq 1.$ We have
$s(A,B)$ $=$ $\frac{1}{4}\mathrm{Tr}\left[\left(A-B\right)^{\ast}\left(A-B\right)+\left(\sqrt{I+A^{2}}-\sqrt{I+B^{2}}\right)^{2}\right]$
$=$ $\frac{1}{4}\left[\left\|A-B\right\|_{2}^{2}+\left\|\sqrt{I+A^{2}}-\sqrt{I+B^{2}}\right\|_{2}^{2}\right].$
The second term here can be evaluated similarly to the second term in (40) leading to
$s(A,B)$$\leq$ $\frac{1}{4}\left[\left\|A-B\right\|_{2}^{2}+\left\|A^{2}-B^{2}\right\|_{1}\right]$
$\leq$ $\frac{1}{4}\left[\left\|A-B\right\|_{2}^{2}+\left(\left\|A\right\|+\left\|B\right\|\right)\left\|A-B\right\|_{1}\right]$
$=$ $\frac{1}{4}\left[\left\|A-B\right\|_{2}^{2}+2\left\|A-B\right\|_{1}\right].$
From the first inequality in (52) we then obtain
$\left[\mathrm{Tr}\sqrt{\rho_{A}}\sqrt{\rho_{B}}\right]^{2}$$\geq$ $\sqrt{1-s(A,B)}\sim 1-\frac{1}{2}s(A,B)$ (53)
$\geq$ $1-\frac{1}{8}\left[\left\|A-B\right\|_{2}^{2}+2\left\|A-B\right\|_{1}\right].$
When both $\rho_{A}$ and $\rho_{B}$ are pure, (51) reduces to
$\mathrm{Tr}\sqrt{\rho_{A}}\sqrt{\rho_{B}}$ $=$ $4]{\det\left(\frac{I-AB}{2}\right)\left(\frac{I-BA}{2}\right)}$
$=$ $4]{\det\left[-\left(\frac{A+B}{2}\right)^{2}\right]}$
$=$ $\sqrt{\det\left(\frac{A+B}{2}\right)},$
also $s(A,B)=\frac{1}{4}\left\|A-B\right\|_{2}^{2}.$ By using the estimates (36), (52) we get
$\left\|\rho_{A}-\rho_{B}\right\|_{1}$ $=$ $2\sqrt{1-\mathrm{Tr}\sqrt{\rho_{A}}\sqrt{\rho_{B}}}$
$=$ $2\sqrt{1-\sqrt{\det\left(\frac{A+B}{2}\right)}}$
$\leq$ $2\sqrt{1-\sqrt[4]{1-\frac{1}{4}\left\|A-B\right\|_{2}^{2}}}$
provided $\left\|A-B\right\|_{2}\leq 2.$ For small $\left\|A-B\right\|_{2}$ this is $\sim\frac{1}{2}\left\|A-B\right\|_{2}.$ The second inequality in (52) implies a lower bound:
$\left\|\rho_{A}-\rho_{B}\right\|_{1}\geq 2\left(1-\exp\left(-\frac{\left\|A-B\right\|_{2}^{2}}{16}\right)\right).$
For further results concerning the trace-norm distance of fermionic Gaussian states see the recent paper [12].